OpenStudy (nwolf):

find the slope (6,3) and (9,-7) would the slope be -7/9 and if so how??

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-7-3)/(9-6) m = -10/3 So the slope of the line that passes through the two points is -10/3

OpenStudy (nwolf):

what points would I graph? would it be the two that was given?

jimthompson5910 (jim_thompson5910):

Yes you would plot the points you were given and draw a line through them.

OpenStudy (anonymous):

m = (y2-y1)/(x2-x1) m = (-7-3)/(9-6) m = -10/3 so slop will be -10/3

jimthompson5910 (jim_thompson5910):

Then you would measure the rise and the run to find the slope

jimthompson5910 (jim_thompson5910):

Alii please don't copy and paste my answer...esp when it's this late...